OpenStudy (anonymous):

A skier has an acceleration of 2.5 m/s2. How long does it take her to come to a complete stop from a speed of 18 m/s? A. 3.4 s B. 7.2 s C. 15.5 s D. 0.14 s

OpenStudy (anonymous):

\[\color{green}{\vec{v}_f = \vec{v}_i + \vec{a}t}\] The final velocity is zero, and you are given the magnitude of the acceleration and the magnitude of the initial velocity. You can assume that the acceleration acts in the opposite direction of the initial velocity, since the object is slowing down.

OpenStudy (anonymous):

the correct answer will be 7.2 because a = v/t so t =v/a